OpenStudy (anonymous):

Simplify: 3radical7/5radical7 A) 7 1/5 B)7 8/15 C)7 5/3 D)7 2/15

OpenStudy (anonymous):

\[\sqrt[3]{7}\div \sqrt[5]{7}\]

OpenStudy (anonymous):

\[\huge 7^{\frac{1}{3}-\frac{1}{5}}\] is a start, i.e. subtract the exponents

jimthompson5910 (jim_thompson5910):

satellite73 is using this rule